RJ is a 55-year-old female patient who is 5 feet 4 inches tall and weighs 112 lb. She is bedridden (activity factor = 1.2) and has a fractured pelvis due to an automobile accident (stress factor 1.4). For this patient, determine (a) TDE, (b) milliliters of 8.5% amino acid solution, and (c) milliliters of 50% dextrose solution.

I already know the answers but I'm not sure how to work it out. Thanks.
Answer #1

Patients name is RJ

Age is 55years

Height of the patient is 5feet 4inches = 162.5cm

Weight - 112lbs = 112×0.45 = 50.4kg

Activity factor - 1.2

Stress factor - 1.4

A. TDE is total daily expenditure .

1. Calculate the total daily expenditure of energy (TDE). TDE = BEE x activity factor x stress factor

= BEE - 66.5+ 13.8( weight in kg ) +5(height in cm) -6.8( age in years )

= 66.5 + 13.8× 50.4 + 5× 162.5- 6.8× 55

= 66.5 + 695.5+ 812.5- 374

= 1200.5 kcal

Now, TDE = 1200.5× 1.2×1.4

= 2016.8

b.) Calculate the volume of amino acids (protein) required.

(mL amino acids = g protein / % concentration amino acid stock solution)

Daily requirement of protein is 0.8g/kg

Weight of the patient is 50.4 kg

= 0.8×50.4 = 40.32g

1 gm of protein provides 4kcal

Then 40.32gm will provide = 4× 40.32

= 161.28 kcal

Now ,ml of amino acid = 40.32 / 8.5%

= 40.32 / 8.5 × 100

= 474.3ml of amino acid /day

c.) Calculate the volume of dextrose (carbohydrate) required.

(mL dextrose = g dextrose / % concentration dextrose stock solution)

= amount of carbohydrate requirements = 362 gm

Now, ml of dextrose = 362/ 50%

= 362 / 50 × 100

= 724 ml dextrose solution /day